OpenStudy (anonymous):

If B is between A and C find the value of BC. AB = 3, BC = 5x - 1, AC = 12 9 8 2 10

OpenStudy (anonymous):

AB = 3, BC = 5x-1, AC = 12 AB + BC = AC 3 + 5x-1 = 12 5x +2 = 12 5x = 12 - 2 5x = 10 x = 2 BC = 5x - 1 = 5*2 - 1 = 10 - 1 = 9

OpenStudy (anonymous):

if B is between A and C ... and given AB=3 and AC=12 then BC=AC-AB=12-3